METCRAX: Logbook Entries

METCRAX: Site Rim Messages, 16 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
150 Sun 29-Oct-2006Rim sonic boom angleRimSemmer
146 Fri 27-Oct-2006rim boom anglesRimOncley
142 Fri 27-Oct-2006Prop at rimRimSemmer
140 Thu 26-Oct-2006PROP change at rimRimSemmer
127 Wed 25-Oct-2006Prop test at RimRimSemmer
124 Tue 24-Oct-20062.8m(?) prop at rimRimSemmer
119 Tue 24-Oct-2006Prop head replaced at rimRimSemmer
118 Tue 24-Oct-2006sonic replacement at rimRimSemmer
115 Sun 22-Oct-2006site visit (mast down)RimOncley
114 Sun 22-Oct-2006testing of PSU 376RimOncley
104 Fri 20-Oct-2006Re-installed rim sonicRimOncley
71 Sat 14-Oct-2006rim sensor serial numbersRimHorst
28 Fri 06-Oct-2006sonic and krypton removed from rimRimMartin
23 Fri 06-Oct-2006rim restartedRimMartin
13 Thu 05-Oct-2006Logger fixed for rimRimMartin
12 Thu 05-Oct-2006Enabled data transmission from rimRimHorst

150: Sonic, Site Rim, Sun 29-Oct-2006 09:22:27 MST, Rim sonic boom angle
The new boom angle was shot for the sonic rim with the datascope. The 
measurement was made by hiking south of the tower along the rim trail until
a proper alignment point was found.
The angle was .5 degree.
Refer to entry 146 for another measurement.

146: Sonic, Site Rim, Fri 27-Oct-2006 16:04:41 MST, rim boom angles
I've just tried the "theodolite-GPS" method of shooting the boom angles:
1. use the theodolite telescope to locate a postion aligned with the boom
2. walk with the GPS towards the tower, watching the heading
3. walk the reverse heading back to the theodolite
4. note the lat/lon of the theodolite.

Due to the heat, I couldn't get a good view of the sonic boom.  I guessed
that, when it disappeared from view along the road east of the base, I was
still had a bit further east to go to be along the boom.  At that point,
the walked heading was 001/181, lat/lon was 35d 00.871'N, 111d 01.753'W.
Thus, this estimate would be 181-1?deg for the boom angle.

I also used this method to shoot the prop angle.  This I was more confident
of since I was closer, and it was later in the day with less heat.
The walked heading was 91/271, lat/lon was 35d 01.774'N, 111d 02.130'W.
(The GPS said its error in both cases was 5m.)
The, this estimate would be 271deg, which is close to perpendicular to 180deg
of the sonic (and amazingly close to cardinal 270deg!)

The lat/lon reported by the station is 35d 01.767'N, 111d 01.749'W.  If the
boom angles were exactly 180 and 270, my prediction of the station lat/lon
would be 35d 01.774'N, 111d 01.753'W, which is off by only 0.007'N, 0.004'W
or ~14m, or off in boom angle by about 1 degree over these distances.  Thus,
I think 181/271 should be used.

142: Prop_Vane, Site Rim, Fri 27-Oct-2006 10:19:16 MST, Prop at rim
The 10m prop at rim does read lower than the 3m sonic so this may be a true
event we have been seeing there.

140: Prop_Vane, Site Rim, Thu 26-Oct-2006 17:05:10 MST, PROP change at rim
The prop has been changed at the rim site. The unit that was at 3m was moved
to the 10m height. The boom direction was changed to 270 degrees at ~17:00
local time..
This work took place between 15:30 and 16:30 local time.

5/23/07, TWH:

Missing data from 16:01 to 16:18 is perhaps when prop was moved up.
127: Prop_Vane, Site Rim, Wed 25-Oct-2006 10:02:23 MST, Prop test at Rim
The prop test at rim was a success. There was very good agreement between the
~3m unit and the sonic. We will move the 3m prop up to the 10m height.
NOTE: It is too windy this afternoon to change lower the PAM mast. Will try to
do it first thing in the morning.

124: Prop_Vane, Site Rim, Tue 24-Oct-2006 16:54:32 MST, 2.8m(?) prop at rim
Since there is not an IOP going on we decided to run a test at rim.
The 10m prop is consistantly reading lower than the 3m sonic so we mounted
another prop at about 2.8m. The prop data now is at the lower height (decided
to use ttyS4 since it was configured correctly). If the lower prop and sonic
agree then we will move the lower prop to the 10m height.
NOTE: the boom angle was not set therefore direction is bad!
This visit took place from 15:30 to 14:15 local time.

119: Prop_Vane, Site Rim, Tue 24-Oct-2006 09:22:12 MST, Prop head replaced at rim
This work took place on Oct 23.
After replacing the sonic at rim, it was noted that the prop speed was
consistently lower than the sonic. When replacing the sonic Steve O. noticed
that the prop head bearings sounded rough. So replaced the prop head.
Mast down approximately between 11:00 and 12:00.

118: Sonic, Site Rim, Tue 24-Oct-2006 09:18:48 MST, sonic replacement at rim
This event took place on Oct 22.
The sonic at rim was replaced with the PSU sonic 376. A serializer was mounted
in this unit since the NCAR sonic from wl failed to work.
This work took place between 12:00 and 13:00.

115: Station_Visit, Site Rim, Sun 22-Oct-2006 13:32:14 MST, site visit (mast down)
Visited rim to:
- exchange sonics (earlier visit by Steve S. retrieved sonic electronics with
  serializer, which was swapped into PSU 376 in the base).
- data download
- shoot boom angles (sonic: 358, prop 108 -- clearly something is wrong --
  probably compass dip angle)
- measure heights of TRH (89.5" from roughly ground to top of boom) and
  GPS (roughly 48" above local ground)
- Prop-vane bearing is a bit "growly", but seemed okay

Mast down about 12:20-12:45.
114: Sonic, Site Rim, Sun 22-Oct-2006 11:04:41 MST, testing of PSU 376
The PSU sonic that we are about to install at rim is reading in a shipping box
with foam cut-out

-1.0 cm/s U
+0.7 cm/s V
-1.0 cm/s W

(values are approximate, read from the CSI plot using +/- 2cm/s full scale)

We think this is good enough!

104: Sonic, Site Rim, Fri 20-Oct-2006 18:28:35 MST, Re-installed rim sonic
We reinstalled the rim sonic (6xx) late this afternoon.  The mast would have
been down from about 16:15-17:15.  We also installed krypton 1394 with it.
(I note that the config now shows 1397 for this site, but 1397 probably is
being used in the crater instead of 1101, which is in the base with a "bad"

Unfortunately, the data for this sonic seems bad despite its working this
morning at wl.0.5m and later testing in the base trailer.  Perhaps it was
damaged as we packed it up without using its (heavy) shipping box, but
Steve and I are both skeptical of this hypothesis.  We're not sure what
action to take to remedy this problem.  The symptom is that it reports some
values, but with a status nibble of "8" for about 5 seconds, then reports
values of all 8000 with a status nibble of "F" for another 5 seconds, then
repeats this cycle.  We assume this is a "reacquire signal" cycle, due to a
weak signal.

TWH:  This is s/n 673; see logbook entry 38.
71: Station_Visit, Site Rim, Sat 14-Oct-2006 12:08:32 MST, rim sensor serial numbers
Q7		Q99259
pygin		EP31977
pygout		EP31980
pyrin		KZ940181
pyrout		KZ940186
Tsoil		STP98032
HFT		H023009
Licor		PY10835
TP01		200234

28: Sonic, Site Rim, Fri 06-Oct-2006 17:09:51 MST, sonic and krypton removed from rim
The sonic, krypton, cabling and brackets were removed from rim and 
brought back to the base, so that they can be installed at sw.
S/N 0673

23: Adam, Site Rim, Fri 06-Oct-2006 11:24:35 MST, rim restarted
rim rebooted. Data collection started at 1820UTC

13: Station_Visit, Site Rim, Thu 05-Oct-2006 13:45:19 MST, Logger fixed for rim
Enabled the data logger for rim this morning. The
power connection from the battery to the campbell 
and radio had been disconnected.

12: Data, Site Rim, Thu 05-Oct-2006 12:57:34 MST, Enabled data transmission from rim
Charlie turned on data transmission  from the rim.